OpenStudy (anonymous):

let vector OP = 2i + 3j and vector OQ = -6i+4j. R is a point on vector PQ such that PR/RQ = k/1, where k>0. a/ express vector OR in terms of k, i and j. b/ express (vector OP)dot(vector OR) and (vector OQ)dot(vector OR) in terms of k. c/ Find the value of k such that OR bisects angle POQ.

OpenStudy (anonymous):

\(OP\bullet OQ=0\)

OpenStudy (anonymous):

thanks what i am having the most trouble with is part c

OpenStudy (anonymous):

 you could use the law of sines...

OpenStudy (anonymous):

\[\frac{PR}{\sin 45}=\frac{OR}{\sin a}\text{ and }\frac{QR}{\sin 45}=\frac{OR}{\sin b}\]\[\Rightarrow \frac{PR}{QR}=\frac{\sin b}{\sin a}\text{ but }a=90-b\Rightarrow \frac{PR}{QR}=\tan b\] \[\cos b = \frac{OQ\bullet QP}{|\!|OQ|\!||\!|QP|\!| }\]

OpenStudy (anonymous):

never thought of using law of sines for this vector problem

OpenStudy (anonymous):

oops. \[\tan b = \frac{|\!|OP|\!|}{|\!|OQ|\!|}\]

OpenStudy (anonymous):

so that's k, right?

OpenStudy (anonymous):

yes its good

OpenStudy (anonymous):

did you draw a pic when you tried this prob?

OpenStudy (anonymous):

yes but i got stuck after doing part b

OpenStudy (anonymous):

what class is this?

OpenStudy (anonymous):

precalculus honors

OpenStudy (anonymous):

doing vectors? well, good on you!

OpenStudy (anonymous):

well some teachers skip vectors but my teacher insists

OpenStudy (anonymous):

that's good. if you ever take physics or linear algebra, you'll be ahead of the game!

OpenStudy (anonymous):

you understand the prob now? I think it's always good to draw pics and label stuff. it can help generate ideas. and don't discard your intuition until you work it through. i've tossed ideas that were valid because i thought they would work but I didn't follow through on them to see the light at the end. i'm sure you know and have a good feel for this. good luck!

OpenStudy (anonymous):

thanks and good night

OpenStudy (anonymous):

thanks for the great question!